OpenStudy (anonymous):

find 2 numbers whose sum is 12 and whose product is the maximum possible value. (hint: let x be one number, then 12-x the other number. Form a quadratic function by multiplying them and then find the maximum value of the function

OpenStudy (science0229):

Can you use a graphing calculator, or is this by hand?

OpenStudy (science0229):

If you're doing this by hand, as the hint states, create a quadratic, which is y=x(12-x).

OpenStudy (anonymous):

does not specifically say, i guess i can use both, but don't know how

OpenStudy (science0229):

Ok.

OpenStudy (science0229):

Did you learn Calculus 1?

OpenStudy (anonymous):

no, i jumped from math 095 to math 103, all this is new to me... thanks for the help

OpenStudy (science0229):

Ok. Then, let's do it the graphing calculator way.

OpenStudy (anonymous):

Isn't the answer just 6 for both numbers?

OpenStudy (science0229):

Yes...

OpenStudy (anonymous):

both numbers

OpenStudy (science0229):

Basically, x=6

OpenStudy (science0229):

Do you know how, though?

OpenStudy (anonymous):

well, actually, i'm trying to figure it out and got 6 and -6 not sure if that's right... how did you get just 6

OpenStudy (science0229):

Do you have a graphing calculator right now?

OpenStudy (anonymous):

oh wait, did you just pick 6 because it asks for the MAXIMUM value of the function

OpenStudy (anonymous):

yes i do

OpenStudy (science0229):

Yes to your first question. If you want to verify it, follow the steps. 1. Graph y=x(12-x) 2. Set your window as the following; domain=from 0 to 12, range=from -10 to 40 3. You should see a vertex, which is your maximum point. 4. Press 2nd+Trace, and press maximum. 5. Set your lower bound and upper bound, so that the vertex is in between. 6. The x-coordinate of the vertex is your x-value that makes the biggest possible product.

OpenStudy (science0229):

That is if you have ti-83 or ti-84

OpenStudy (anonymous):

yes ti-84

OpenStudy (science0229):

Then, follow the steps to get x=6.

OpenStudy (anonymous):

sorry, don't know how to figure it out on calculator either

OpenStudy (anonymous):

\[ y=x(12-x)= -x^2+12 x = -x^2 +12 x -36 +36= -(x-6)^2+36 \] This a parabola of vertex (6,36} turning down. The maximum is when both numbers are equal to 6. 6+6 =12 6 x 6 =36

OpenStudy (anonymous):

where did you get the 36?

OpenStudy (anonymous):

just to complete the square.

OpenStudy (anonymous):

I added 36 and subtracted 36 so I added 0

OpenStudy (anonymous):

Have you studied parabolas?

OpenStudy (anonymous):

yes i have, a little confused about it, i but i think i got it

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

YW